can you determine the type and the solution set of this system of equation:
5x-2-5=0
3x+y-3=0 ?

Found 2 solutions by mananth, richwmiller:
Answer by mananth(16946)


5x-2y-5=0
3x+y-3=0 ?
5x-2=5
3x+y=3
compare with equations a1x+b1y+c1=0 & a2x+b2y+c2=0
a1/a2= 5/3
b1/b2= -2
c1/c2= 5/3
a1/a2 not equal to b1/b2
so unique solution

Answer by richwmiller(17219)

Did you copy the first equation correctly? There is no "y" !
5x-2y-5=0
3x+y-3=0

5x-2y=5
3x+y=3

5,-2,5
3,1,3
One solution (1,0)